OpenStudy (anonymous):

What is the perimeter of a room,6.2m length and 4m breath ?

OpenStudy (ankitshaw):

room is rectangle in shape so two time the length + two times the breadth = perimeter so 2x6.2 + 2x4= 12.4+8 = 20.4

OpenStudy (ankitshaw):